A company manufactures three types of t.vs A,B and C. The company manufactures four less type B than twice A, and one more type B than the sum of types A and type C. If the total number of t.vs manufactured per week is 51. How many of each type of t.v are manufactured.
Can someone please help this question has me in a twist!

Found 2 solutions by mathstutor494, kayroge:
Answer by mathstutor494(120)

Let type A be x
Then type B = 2x+4
And B-1 = x+C
Substituting for B
2x+4-1 = x+C
So C = x+3
Given that total number of tvs manufactured per week is 51 which is A+B+C
so
x+ (2x+4) + (x+3) = 51
4x+7 = 51
4x = 44 so x =11
Type B = 2x+4 = 2*11+4 = 26
and C = x+3 = 11+3 = 14
Ans Type A= 11, Type B = 26 & type C = 14

Answer by kayroge(2)

So let A be A, B be B and C be C
2A-4=B
and
2A-4+1=A+C
2A-A-4+1=C
so A-5=C
so
A+2A-4+A-5=51
4A-4-5=51
4A=51+4+5
A=60/4
A=15
So B=B+15x2-4=26
So C=51-26-15=10
Answer equals A=15 B=26 and C=10